OpenStudy (anonymous):

Use f(x) = 2x − 5 and g(x) = 3 − x2 to evaluate the expression (g o g)(x)

jimthompson5910 (jim_thompson5910):

\[\large g(x) = 3-x^2\] \[\large g(g(x)) = 3-(g(x))^2\] \[\large g(g(x)) = 3-(3-x^2)^2\] \[\large g(g(x)) = 3-(3-x^2)(3-x^2)\] \[\large g(g(x)) = 3-(9-3x^2-3x^2+x^4)\] \[\large g(g(x)) = 3-(9-6x^2+x^4)\] \[\large g(g(x)) = 3-9+6x^2-x^4\] \[\large g(g(x)) = -x^4+6x^2-6\]